OpenStudy (anonymous):

convert to rectangular form r = 4 / (7sin theta + 29 cos theta)

terenzreignz (terenzreignz):

Divide both sides by r. What do you get?

OpenStudy (anonymous):

honestly thats what I have been having an issue with... I can't figure out how to switch the r because of the fraction

terenzreignz (terenzreignz):

Yeah, well, if you divide both sides by r, you should get \[\Large 1 = \frac{4}{r[7\sin(\theta)+29\cos(\theta)]}\] right?

terenzreignz (terenzreignz):

Might have been an inefficient approach.... multiply both sides here by \[\large 7\sin(\theta) + 29 \cos(\theta) \] instead. Giving you... \[\Large r[7\sin(\theta) + 29 \cos (\theta)]=4\]

OpenStudy (anonymous):

ah okay I got it now I didnt even think to multiply both sides by the denominator now its just 7y + 29x = 4

OpenStudy (anonymous):

and solve for y for the answer correct?

terenzreignz (terenzreignz):

Honestly, I think 7y+29x = 4 is enough But solving for y is good too :)

OpenStudy (anonymous):

awesome thank you very much